Which of the following is a characteristic of pure monopoly?

Question 2

Describe the characteristics of the long-run equilibrium for a perfectly competitive firm.

Assume a purely competitive firm is selling 200 units of output at $3 each. At this output, its total fixed cost is $100 and its total variable cost is $350. This firm

is making a profit, but not necessarily the maximum profit.

is maximizing its profit.

is incurring losses.

should shut down in the short run.

Assume a purely competitive firm is selling 200 units of output at $3 each. At this output, its total fixed cost is $100 and its total variable cost is $350. This firm

is making a profit, but not necessarily the maximum profit.

is maximizing its profit.

is incurring losses.

should shut down in the short run.
Answer #1

Answer : 1) Characteristics of pure monopoly are following :

a) There exists only one seller in the market.

b) In pure monopoly the seller has market power to set the price level.

c) There exists a high entry barrier in pure monopoly market.

2) The answer is option A.

Total cost = Total fixed cost + Total variable cost = 100 + 350 = $450

Average total cost = Total cost / Quantity = 450 / 200 = $2.25

The price is $3 per unit (Given).

As here the price is higher than the average total cost hence the firm earns profit but not necessarily maximum profit. Hence except option A other options are not correct. Therefore, option A is the correct answer.